A seminar is being scheduled on high school mathematics. There are 5 speakers in attendance: Mr. Bower, Mr.Smith, Ms. Patel, Ms. Volpe and Ms. Berry. What is the probability that:

a. They will speak in alphabetical order?
b. Ms. Patel speaks first so she can leave early?

c. The women will not speak one after another?
Answer #1

Number of speakers = 5

Number of orders possible = 5! = 120

a. P(they will speak in alphabetical order) = 1/120

b. P(Ms. Patel speaks first) = 1/5

c. P(the women will not speak one after other) = P(first is woman, second is man, third is woman, 4th is man and 5th is woman)

= 3/5 x 2/4 x 2/3 x 1/2 x 1/1

= 1/10